OpenStudy (anonymous):

A drill began digging a well 48 3/5 feet above sea level.When the work was complete,the drill had dug in to the ground to a depth of 12 1/2 feet below sea level.How far did the drill dig in all?

OpenStudy (anonymous):

so final height - initial height = distance. Think sea level as 0, so -12 1/2 - 48 3/5 = ?

OpenStudy (anonymous):

and then take the absolute value of the answer, because distance can't be negative.

OpenStudy (anonymous):

Wait...wouldn't it be 48 3/5 - -12 1/2?

OpenStudy (anonymous):

same either way.

OpenStudy (anonymous):

I still don't get it

OpenStudy (anonymous):

Why subtract?

OpenStudy (anonymous):

It's the total distance

OpenStudy (anonymous):

\[\left(48+\frac{3}{5}\right)-\left(-12-\frac{1}{2}\right)=\frac{174}{5}=34+\frac{4}{5} \]

OpenStudy (anonymous):

vicky191, think about it this way. if a plane starts 2m above ground and stops at 4m above ground, Wouldn't you do 4-2 to find the distance? ( you could also do 2-4 and take the absolute value. Either way the answer will be 2m.) Similarly, you could do -12 1/2 - 48 3/5 as well as 48 3/5 - (-12 1/2).